# Tracezip: Efficient Distributed Tracing via Trace Compression

### v1

@@ 3.3. Span Retrieval Compression and Uncompression
21:     Compress span with the corresponding path identifier and the values of local fields
@@
Our idea is that for spans generated in each service instance, we organize their key-value pairs as a prefix-tree-like data structure, referred to as Span Retrieval Tree (SRT). The SRT functions as a multi-way tree, with all non-leaf nodes (except for the root) associated with a key-value pair. For each type of span, identified by a unique span Name, there is only one leaf node connected to all the last non-leaf nodes stemming from it. This leaf node holds a collection of keys without values. Figure 4 illustrates an example of SRT, where the gray node and the yellow nodes represent the root and the leaves, respectively, while the remaining are the non-leaf nodes. Each span can be “spelled out” by tracing a path from the root down to the leaf. The path of SRT represents the set of KV pairs shared across multiple spans. The non-leaf nodes contain the fields that are more repetitive, which we refer to as universal fields. Although spans may exhibit commonality, they will still have some unique KV pairs, such as those related to ID and time information. We refer to such pairs as local fields and only store their keys at the leaf. The rationale is that such unique fields are incompressible, i.e., not shared with other spans, so we discard their values. Based on SRT, a span can be represented as a unique path identifier plus its exclusive values that are extracted according to the keys in the leaf node. Each path identifier collectively represents the KV pairs shared among spans, instead of one identifier for each key and value. Since these common KV pairs constitute a significant portion, the trace size can be substantially reduced, enhancing the overall efficiency.
@@ 3.4.2. Mapping-based Tree Compression
Although we have restructured the SRT to eliminate redundant nodes, there could still be repeated keys and values in it. For example, in Figure 4, key data.size appears in all data.size nodes, e.g., {“data.size”: “64 types”} and two of them also share value 64 types. Thus, to further compress the size of SRT, we employ a dictionary to map keys/values that occur multiple times to shorter identifiers. We construct the identifiers using the standard alphanumeric set, i.e., [0-9a-zA-Z]. Initially, the hashed output consists of a single character, from ’0’ to ’9,’ followed by ’a’ through ’z,’ and finally ’A’ through ’Z.’ Upon exhausting the single character possibilities, the function increases the length of the hash output to two characters, starting from ’01,’ and so forth. Since each universal field has limited distinct values, i.e., smaller than $\psi$, the dictionary will also be small in size. Note we do not encode the values of local fields (not in the SRT), which may inevitably make the dictionary too big given their diversity. Similar to the SRT synchronization process between services and tracing backend, the dictionary will be sent to the backend every time it undergoes an update.
@@ 5.1.1. Deployed Cloud Systems
Train Ticket is a railway ticketing application comprising 41 microservices, each responsible for a specific function, such as user authentication, ticket booking, payment processing, and notification. This benchmark is implemented in different programming languages such as Java, Go, Node.js, Python, etc. Train Ticket allows a comprehensive evaluation in a multifunctional scenario, which has been widely used in many trace-related topics, including trace sampling (Chen et al., 2024), root cause localization (Zhou et al., 2021, 2023), service architecture measurement (Peng et al., 2022), etc. In order to replicate a live production environment, we employed Locust (Locust, 2024), an open-loop asynchronous workload generator, to drive the services. The workloads are directly borrowed from the original work (Zhou et al., 2021) that introduces the Train Ticket microservices.
@@ 5.2.1. Open-source Cloud Systems
Table 5.1.3 presents the compression performance when collecting traces of the microservices benchmark and cloud applications components. For each system, we calculate the total size of traces collected, the size after compression, and the resultant compression ratios (CRs) when applying different compression algorithms. We can see that Tracezip, as a standalone solution, can achieve CRs ranging from 3.94 to 6.46. This demonstrates that Tracezip can remove more amount of redundant information than that shown by our preliminary study in Section 2.2. Traditional compression tools, i.e., gzip, bzip2, and lzma, reduce the file size with a combination of different techniques such as dictionary-based compression and Huffman coding. Among them, bzip2 generally outperforms the others across most systems, with gzip having the least effectiveness. In the Train Ticket benchmark, the tools demonstrate the least effective compression with a CR of roughly 10, while on the cloud application components, they deliver a better performance, attaining a comparable CR of about 20.
@@ 5.3.1. Space Complexity
At the service side, Tracezip maintains three types of data structures to capture the redundancy among trace spans and perform compression, namely, a SRT along with its hashed paths and a map for dictionary-based compression. To prevent impeding the normal execution of the service, it is imperative that they are constrained in size without excessive memory consumption. To study the space complexity of Tracezip, we select a microservice in Alibaba trace data with diverse spans and calculate the cumulative size of the three data structures after the compression procedure. A critical parameter influencing this size is the threshold $\psi$, which dictates the maximum number of distinct values for universal attributes. An attribute having an exceeding number of values will be moved to the leaf node, becoming a local attribute. A larger $\psi$ enables Tracezip to compress a broader spectrum of span fields, enhancing performance but at the cost of a more substantial SRT and mapping structure. Conversely, a small $\psi$ compromises the effectiveness but with lower space overhead in return.
Figure 7. Performance with Different $\psi$
@@ 5.4. Threats to Validity
Internal validity. One of the primary concerns regarding internal validity is the accuracy of our evaluation metrics and the potential biases in our experimental setup. Real-world cloud services exhibit a vast array of complexities and variations, making it challenging to capture all possible scenarios within a single study. To address this challenge, we carefully selected a diverse set of microservices benchmarks and production trace data from Alibaba that we believe are representative of typical cloud service operations. These datasets were chosen to reflect common patterns and behaviors observed in real-world applications, thus providing a meaningful context for evaluating Tracezip’s performance. Additionally, any configuration or tuning of Tracezip that is specific to these datasets might inadvertently favor our approach, potentially skewing the results. To mitigate this, we ensured that the benchmarks and datasets were selected and configured independently of Tracezip’s development process.
External validity. External validity pertains to the generalizability of our results to other settings or systems. Our evaluation of Tracezip was specifically designed to address the diversity inherent in real-world cloud systems. We implemented and tested Tracezip within the OpenTelemetry Collector framework and evaluated it across a range of cloud environments and backend components, including gRPC, Apache Kafka, MySQL, and others. These settings were carefully selected to reflect the variety of systems and technologies commonly used in cloud services, ensuring a comprehensive basis for assessing Tracezip’s effectiveness. By choosing such a diverse array of environments and applications, we aimed to capture the broad spectrum of redundancy patterns and data characteristics found in typical cloud systems. This approach helps to ensure that our findings are applicable to a wide range of real-world scenarios, demonstrating Tracezip’s capability to perform effectively in diverse and dynamic cloud environments.
6. Discussion
@@ 7.1. Distributed Tracing Systems
Distributed tracing offers a holistic, end-to-end perspective on data requests as they navigate the myriad services within a distributed application. Pioneering works, such as Magpie (Barham et al., 2003), Whodunit (Chanda et al., 2007), X-Trace (Fonseca et al., 2007), and Dapper (Sigelman et al., 2010), have laid the foundation of tracing in distributed systems. Recent studies take steps further to address specific challenges to pursue accurate and efficient tracing capabilities. Pivot Tracing (Mace et al., 2015) and Canopy (Kaldor et al., 2017) emphasize cross-application and cross-platform tracing to ensure seamless monitoring and diagnostics across diverse environments. Panorama (Huang et al., 2018) achieves more sophisticated observability by collecting inter-process and inter-thread error signals. DeepFlow (Shen et al., 2023) is a non-intrusive distributed tracing framework for troubleshooting microservices. It establishes a network-centric tracing plane with eBPF technique in the kernel. OpenTelemetry (OpenTelemetry, 2024a) standardizes the APIs for telemetry data collection, instrumentation libraries, and semantic conventions. It can be used with a broad variety of observability backends, such as Jaeger (Jaeger, 2024), Zipkin (Zipkin, 2024), and Prometheus (Prometheus, 2024).
@@ 8. Conclusion
In this paper, we propose Tracezip, an online and scalable solution to mitigate the computational and storage overhead associated with distributed tracing. Existing trace sampling techniques either sacrifices the accurate detection of edge cases or tracing scalability. As an orthogonal approach, Tracezip reduces tracing overhead by compressing trace spans into a concise representation, substantially reducing the costs of trace transmission and storage. This is achieved by removing the redundant data shared across multiple spans. The full trace spans can then be seamlessly restored by retrieving the common data already delivered by previous spans. At the service side, such redundancy is continuously encapsulated by Span Retrieval Tree (SRT), which will be synchronized at the backend to ensure consistent trace compression and uncompression. To manage the complexity of SRT, Tracezip constantly restructures SRT into its optimal form and employs mapping technique to further reduces its size. Tracezip also encompasses a differential update mechanism for efficiently synchronize SRT between services and the backend. We have implemented Tracezip insight OpenTelemetry Collector and evaluated it on open-source cloud services and production trace data. Experimental results highlight its potential to pave the way for efficient system monitoring.
Acknowledgment

### v2

@@ 3.3. Span Retrieval Compression and Uncompression
21:     Compress span based on the corresponding path identifier and the values of local fields
@@
Our idea is that for spans generated in each service instance, we organize their key-value pairs as a prefix-tree-like data structure, i.e., SRT. The SRT functions as a multi-way tree, with all non-leaf nodes (except for the root) associated with a key-value pair. For each type of span (which is identified by a unique span Name), there is only one leaf node connected to all the last non-leaf nodes stemming from it. This leaf node holds a collection of keys without values. Figure 4 illustrates an example of SRT, where the gray node and the yellow nodes represent the root and the leaves, respectively, while the remaining are the non-leaf nodes. Each span can be “spelled out” by tracing a path from the root down to the leaf. The path of SRT represents the set of KV pairs shared across multiple spans. The non-leaf nodes contain the fields that are more repetitive, which we refer to as universal fields. Although spans may exhibit commonality, they will still have some unique KV pairs, such as those related to ID and timestamps. We refer to such pairs as local fields and only store their keys at the leaf. The rationale is that such unique fields are incompressible, i.e., not shared with other spans, so we discard their values. Based on SRT, a span can be represented as a unique path identifier plus its exclusive values that are extracted according to the keys in the leaf node. Each path identifier collectively represents the KV pairs shared among spans, instead of one identifier for each key and value. Since these common KV pairs constitute a significant portion, the trace size can be substantially reduced, enhancing the overall efficiency.
@@ 3.4.2. Mapping-based Tree Compression
Although we have restructured the SRT to eliminate redundant nodes, there could still be repeated keys and values in it. For example, in Figure 4, key data.size appears in all data.size nodes, e.g., {“data.size”: “64 bytes”} and two of them also share value 64 bytes. Thus, to further compress the size of SRT, we employ a dictionary to map keys/values that occur multiple times to shorter identifiers. We construct the identifiers using the standard alphanumeric set, i.e., [0-9a-zA-Z]. Initially, the hashed output consists of a single character, from ’0’ to ’9,’ followed by ’a’ through ’z,’ and finally ’A’ through ’Z.’ Upon exhausting the single character possibilities, the function increases the length of the hash output to two characters, starting from ’00,’ ’01,’ and so forth. Since each universal field has limited distinct values, i.e., smaller than $\psi$, the dictionary will also be small in size. Note we do not encode the values of local fields (not in the SRT), which may inevitably make the dictionary too big given their diversity. Similar to the SRT synchronization process between services and the tracing backend, the dictionary will be sent to the backend every time it undergoes an update.
@@ 5.1.1. Deployed Cloud Systems
Train Ticket is a railway ticketing application comprising 41 microservices, each responsible for a specific function, such as user authentication, ticket booking, payment processing, and notification. This benchmark is implemented in different programming languages such as Java, Go, Node.js, Python, etc. Train Ticket allows a comprehensive evaluation in a multifunctional scenario, which has been widely used in many trace-related topics, including trace sampling (Chen et al., 2024), root cause localization (Zhou et al., 2021, 2023), service architecture measurement (Peng et al., 2022), etc. In order to replicate a live production environment, we employ Locust (Locust, 2024), an open-loop asynchronous workload generator, to drive the services. The workloads are directly borrowed from the original work (Zhou et al., 2021) that introduces the Train Ticket microservices.
@@ 5.2.1. Open-source Cloud Systems
Table 5.1.3 presents the compression performance when collecting traces of the microservices benchmark and cloud applications components. For each system, we calculate the total size of traces collected, the size after compression, and the resultant compression ratios (CRs) when applying different compression algorithms. We can see that Tracezip, as a standalone solution, can achieve CRs ranging from 3.94 to 6.46. This demonstrates that Tracezip can remove more amount of redundant information than that shown by our preliminary study in Section 2.2. Traditional compression tools, i.e., gzip, bzip2, and lzma, reduce the file size with a combination of different techniques such as dictionary-based compression and Huffman coding. Among them, bzip2 generally outperforms the others across most systems, with gzip having the least effectiveness. In the Train Ticket benchmark, the tools demonstrate the least effective compression with a CR of roughly 10, while on the cloud application components, they deliver a better performance, attaining a comparable CR of around 20.
@@ 5.3.1. Space Complexity
At the service side, Tracezip maintains two types of data structures to capture the redundancy among trace spans and perform compression, namely, a SRT along with its hashed paths and a map for dictionary-based compression. To prevent impeding the normal execution of the service, it is imperative that they are constrained in size without excessive memory consumption. To study the space complexity of Tracezip, we select a microservice in Alibaba trace data with diverse spans and calculate the cumulative size of the three data structures after the compression procedure. A critical parameter influencing this size is the threshold $\psi$, which dictates the maximum number of distinct values for universal attributes. An attribute having an exceeding number of values will be moved to the leaf node, becoming a local attribute. A larger $\psi$ enables Tracezip to compress a broader spectrum of span fields, enhancing performance but at the cost of a more substantial SRT and mapping structure. Conversely, a small $\psi$ compromises the effectiveness but with lower space overhead in return.
Figure 6. Compression on Alibaba Production Traces
@@ 5.4. Threats to Validity
Internal validity. One of the primary concerns regarding internal validity is the accuracy of our evaluation metrics and the potential biases in our experimental setup. Real-world cloud services exhibit a vast array of complexities and variations, making it challenging to capture all possible scenarios within a single study. To address this challenge, we carefully select a diverse set of microservices benchmarks and production trace data from Alibaba that we believe are representative of typical cloud service operations. These benchmarks and dataset are chosen to reflect common patterns and behaviors observed in real-world applications, thus providing a meaningful context for evaluating Tracezip’s performance. Additionally, any configuration or tuning of Tracezip that is specific to these datasets might inadvertently favor our approach, potentially skewing the results. To mitigate this, we ensure that the benchmarks and dataset are selected and configured independently of Tracezip’s development process.
External validity. External validity pertains to the generalizability of our results to other settings or systems. Our evaluation of Tracezip is specifically designed to address the diversity inherent in real-world cloud systems. We implement and test Tracezip within the OpenTelemetry Collector framework and evaluate it across a range of cloud environments and backend components, including gRPC, Apache Kafka, MySQL, and others. These settings are carefully selected to reflect the variety of systems and technologies commonly used in cloud services, ensuring a comprehensive basis for assessing Tracezip’s effectiveness. By choosing such a diverse array of environments and applications, we aim to capture the broad spectrum of redundancy patterns and data characteristics found in typical cloud systems. This approach helps to ensure that our findings are applicable to a wide range of real-world scenarios, demonstrating Tracezip’s capability to perform effectively in diverse and dynamic cloud environments.
@@ 6.1. Distributed Tracing Systems
Distributed tracing offers a holistic, end-to-end perspective on data requests as they navigate the myriad services within a distributed application. Pioneering works, such as Magpie (Barham et al., 2003), Whodunit (Chanda et al., 2007), X-Trace (Fonseca et al., 2007), and Dapper (Sigelman et al., 2010), have laid the foundation of tracing in distributed systems. Recent studies take steps further to address specific challenges to pursue accurate and efficient tracing capabilities. For example, Pivot Tracing (Mace et al., 2015) and Canopy (Kaldor et al., 2017) emphasize cross-application and cross-platform tracing to ensure seamless monitoring and diagnostics across diverse environments. Panorama (Huang et al., 2018) achieves more sophisticated observability by collecting inter-process and inter-thread error signals. DeepFlow (Shen et al., 2023) is a non-intrusive distributed tracing framework for troubleshooting microservices. It establishes a network-centric tracing plane with eBPF technique in the kernel. OpenTelemetry (OpenTelemetry, 2024a) standardizes the APIs for telemetry data collection, instrumentation libraries, and semantic conventions. It can be used with a broad variety of observability backends, such as Jaeger (Jaeger, 2024), Zipkin (Zipkin, 2024), and Prometheus (Prometheus, 2024).
@@ 7. Conclusion
In this paper, we propose Tracezip, an online and scalable solution to mitigate the computational and storage overhead associated with distributed tracing. Existing trace sampling techniques either sacrifices the accurate detection of edge cases or tracing scalability. As an orthogonal approach, Tracezip reduces tracing overhead by compressing trace spans into a concise representation, substantially reducing the costs of trace transmission and storage. This is achieved by removing the redundant data shared across multiple spans. The full trace spans can then be seamlessly restored by retrieving the common data that are already delivered by previous spans. At the service side, such redundancy is continuously encapsulated by Span Retrieval Tree (SRT), which will be synchronized with the backend to ensure consistent trace compression and uncompression. To manage the complexity of SRT, Tracezip constantly restructures SRT into its optimal form and employs mapping technique to further reduces its size. Tracezip also encompasses a differential update mechanism for efficiently synchronize SRT between services and the backend. We have implemented Tracezip within OpenTelemetry Collector and evaluated it on open-source cloud services and production trace data. Experimental results highlight its potential to pave the way for efficient system monitoring.
Acknowledgments